OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 8 m^3. The length of its base is twice the width. Material for the base costs 10 dollars per square meter. Material for the sides costs 4 dollars per square meter. Find the cost of materials for the cheapest such container. Let ell denote the length of the base, w the width of the base, and let h denote the height of the container. First, write an equation for the cost of the materials in terms of ell,w, and h. Cost =

OpenStudy (anonymous):

Now use the information given to solve for ell in terms of w, and to solve for h in terms of w. Using these, we may express the cost as a function of w alone: C(w) =

OpenStudy (anonymous):

Using this, we find the width of the cheapest such container is: and the cost of materials for the cheapest such container is:

OpenStudy (tkhunny):

L = Length of container W = Width of container H = Height of container There. You're half way there! What's next?

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

V= L * W * h

OpenStudy (tkhunny):

Excellent. V = L*W*H = 8 m^3 Now this part: "The length of its base is twice the width"

OpenStudy (anonymous):

2w*w*h?

OpenStudy (anonymous):

h=(8m^3)/(2w^2)

OpenStudy (tkhunny):

Well, I would be happier if you: 1) Explicitly showed L = 2W, and 2) Substituted without changing varible names: V = 2W*W*H = 8 m^3 3) And even happier if you then simplified: V = 2H*W^2 = 8 m^3

OpenStudy (anonymous):

I see, i have no idea what the next step is

OpenStudy (tkhunny):

Now, we're down to the cost of the container. Area of Base? Lateral Surface Area?

OpenStudy (anonymous):

how do i find the cost of the container?

OpenStudy (tkhunny):

Answer the two questions I just posed. Express both areas in terms of W.

OpenStudy (anonymous):

i don't know what i need to don :(

OpenStudy (tkhunny):

Yes you do. Base = L*W Lateral Surface Area = 2*L*H + 2*W*H What about that do you know know how to do?

OpenStudy (anonymous):

oh i see

OpenStudy (tkhunny):

Now, write them in terms of W.

OpenStudy (tkhunny):

L = 2W H=(8m^3)/(2W^2)

OpenStudy (anonymous):

W=Base/H W=-L

OpenStudy (anonymous):

Can you just give me the answer then i can just look through it please?

OpenStudy (anonymous):

then i can find out how i can do this problem

OpenStudy (tkhunny):

No. You are so close. Why are you giving up? Base = L*W Given L = 2W, we have Base = (2W)*W = 2W^2 Are you SURE that's hard? You do the Lateral Surface Area. Lateral Surface Area = 2*L*H + 2*W*H Given L = 2W, and H=(8m^3)/(2W^2) Lateral Surface Area = 2*(_____)*(_____) + 2*W*(_____)

OpenStudy (anonymous):

=2(2w)*((8m^3)/(2W^2))+2*W((8m^3)/(2W^2))

OpenStudy (anonymous):

=16m^3/w+8m^3/2W

OpenStudy (tkhunny):

Excellent. One more thing. That's just the area. We still don't have the Cost. Here's the expression to minimize: 10*(Base Area) + 4*(Lateral Surface Area)

OpenStudy (anonymous):

minimize=20w^2+(96m^3)/w

OpenStudy (anonymous):

W=(-96m^3)/(2w^2)

OpenStudy (tkhunny):

96?

OpenStudy (anonymous):

=-48m^3/w^2

OpenStudy (anonymous):

64+32

OpenStudy (anonymous):

10(2W^2)+4(16m^3/w+8m^3/w)

OpenStudy (anonymous):

20W^2+(64m^3/w+32m^3/w)

OpenStudy (tkhunny):

Lateral Surface Area 16m^3/W+8m^3/2W = 32m^3/2W+8m^3/2W = 40m^3/2W = 20m^3/W

OpenStudy (anonymous):

it is not 2W, it is just W

OpenStudy (anonymous):

i wrote it wrong

OpenStudy (anonymous):

24m^3/w

OpenStudy (tkhunny):

Given L = 2W, and H=(8m^3)/(2W^2) Lateral Surface Area = 16/W + 8/W = 24/W Okay, I'll bite. Be neater and more careful.

OpenStudy (anonymous):

now what is the next step i need to do?

OpenStudy (tkhunny):

You already had it: minimize=20w^2+(96m^3)/w

OpenStudy (anonymous):

do i add them together?

OpenStudy (tkhunny):

Do you want the total cost? You can build a container with only a bottom, can you? Only sides? Yes, add them.